OpenStudy (anonymous):

Points A(1,2), B(2,3), and C (3,6) lie in the coordinate plane. What is the ratio of the slope of line AB to the slope of line AC?

OpenStudy (anonymous):

AB slope =1 3-2=1 2-1=1 1/1=1 AC slope =2 6-2=4 3-1=2 4/2=2 It is a 1:2 ratio